OpenStudy (anonymous):

Okay so this is a simple problem but the decimals are confusing me the problem is x+4.25+3.5x-1.5x-0.75 i have a answer key and its suppose to be 5 but i've gotten 1.6repeating twice

OpenStudy (anonymous):

x+(3.5-1.5)x+(4.25-0.75) =x+2.0x+3.50 =3x+3.5

OpenStudy (anonymous):

umm with that x equals -1.6 repeating.

OpenStudy (anonymous):

Is this an equation ? I mean =0

OpenStudy (anonymous):

it is an equation but it is not equal to zero

OpenStudy (anonymous):

give the complete equation.

OpenStudy (anonymous):

oh opps x+4.25=3.5x-1.5x-0.75 not x+4.25+3.5x-1.5x-0.75

OpenStudy (anonymous):

x + 4.25 + 3.5x - 1.5x - 0.75 you cannot solve for x unless there is an equal sign...this is just an expression not an equation

OpenStudy (anonymous):

x=(3.5-1.5)x-4.25-0.75 x-2x=-5 -x=-5 x=5

OpenStudy (anonymous):

x + 4.25 = 3.5x - 1.5x - 0.75 --- multiply entire equation by 100 100x + 425 = 350x - 150x - 75 -- combine like terms 100x + 425 = 200x - 75 -- subtract 200x from both sides 100x - 200x + 425 = -75 -- subtract 425 from both sides 100x - 200x = -75 - 425 -- combine like terms -100x = -500 x = -500/-100 x = 5

OpenStudy (anonymous):

what i've been doing is adding the x's together on the right side wich makes 2x and then moving all xs to one side which is 3x so the you have 4.25+3x-0.75 and you move the '-0.75' and hae 5=3x and then 5 / 3 = 1.66666....

OpenStudy (anonymous):

you can get rid of the decimal by multiplying the entire equation by 100

OpenStudy (anonymous):

for any problem with a decimal .?

OpenStudy (anonymous):

well....if the decimal is 4.6, then you could multiply by 10...but if it is 4.26, then you have to multiply by 100. It just depends on where the decimal falls

OpenStudy (anonymous):

It can be like this. 4.25+0.75=3.5x-1.5x-1x 5.00=3.5x-(1.5+1.0)x=3.5x-2.5x=1.0 x x=5

OpenStudy (anonymous):

I do not like fractions or decimal in my equations, so I try to get rid of them. It makes solving the equation so much easier

OpenStudy (anonymous):

okay i tried to solve it after mutiplying the entier thing by 100 but still ended up with 1.6 is it because i moved my xs to the opposite side

OpenStudy (anonymous):

1.666...*

OpenStudy (anonymous):

you want your x's on one side and everything else on the other side. Look at the way I did it, and see what you did differently

OpenStudy (anonymous):

oh okay i found what i did wrong it was a simple computation mistake .(: thank you

OpenStudy (anonymous):

glad to help :)